The probability that a student passing a CAT is ¾, if he pass the CAT, the probability of getting grade A finally is 5/6. If he fails the probability of getting on A would be 1/5. If a student gets a grade A what is the probability that he passed the CAT
1
Expert's answer
2020-12-17T18:40:56-0500

The probability P1 of the event that a student passes CAT and gets a grade A is:

P1= 3/4*5/6 = 5/8

The probability P2 of the event that a student does not pass CAT and gets a grade A is:

P2 = 1/4*1/5 = 1/20

The probability P3 of the event that a student gets a grade A is:

P3 = P1+ P2 = 5/8 + 1/20 = 27/40

The probability P4 of the event that a student passed CAT if he gets a grade A is:

P4 = P1 / P3= (5/8) / (27/40) = 25/27